# Free energy in spin glass models with conventional order

### v1

@@ 3. Approach via constrained free energy
We remark that, to our best knowledge, this proposition is new because the convexity of $\xi$ is only assumed to be over $\mathbf{S}^{D}_{+}$ as in (H4). Under a stronger assumption that $\xi$ is convex on $\mathbb{R}^{D\times D}$, this proposition is direct consequence of [24, Theorem 2 and Lemma 2]. The stronger convexity is needed in the proof of the upper bound [24, Lemma 2] to use Guerra’s interpolation.
@@
It is proven in [24, Theorem 2] that
@@
Since $\inf_{\pi:\>\pi(1)=z}\mathscr{P}^{\mathbf{s}}(\pi,y)\geqslant\inf_{\pi}\mathscr{P}^{\mathbf{s}}(\pi,y)=\mathscr{P}^{\mathbf{s}}(y)$, we get the lower bound for (3.6). The Parisi functional in [24, Theorem 2] is written in a different notation. It is explained in [6, Appendix A] how to rewrite it in our notation.
@@
[…]
@@
[…]
@@
The proof for the upper bound is contained in the proof of [24, Lemma 3]. For completeness, we present it below. Temporarily fix any $\delta>0$. By Proposition 3.2 and (3.7), for every $z\in\mathcal{K}$, there is $\varepsilon_{z}>0$ such that
@@
The standard Gaussian concentration inequalities (e.g. [20, Theorem 1.2]) gives a constant $C$ such that $\mathbb{E}\left|\widetilde{F}^{G}_{N}(S)-F^{G}_{N}(S)\right|\leqslant CN^{-\frac{1}{2}}$ uniformly in $N$ and $S$. Therefore, we get
@@ Appendix A Convergence of $m_{N}$
As mentioned in Remark 1.4, we can show that $m_{N}$ in (1.4) always converges under the Gibbs measure associated with $F^{\mathsf{soc}}_{N}$ in (1.2). When $h=\mathbf{s}$, we have that $m_{N}=\frac{\sigma\sigma^{\intercal}}{N}$ is the self-overlap and such a result has been proved in [6, Theorem 1.1 (1) and (2)]. A straightforward modification gives the desired result below.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Recall ~$\widetilde{\mathcal{F}}_{N}$ below (2.6) and we write ~$\widetilde{\mathcal{F}}_{N}(\cdot)=\widetilde{\mathcal{F}}_{N}(0,\cdot)$ for brevity. We can use (2.11) to rewrite
@@
In view of (2.12), ~$\widetilde{\mathcal{F}}_{N}$ is convex, which implies that $y\cdot\nabla\widetilde{\mathcal{F}}_{N}(0)-y\cdot\nabla\mathcal{F}_{N}(0)$ is bounded from above by
~$\frac{\widetilde{\mathcal{F}}_{N}(ry)-\widetilde{\mathcal{F}}_{N}(0)}{r}-y\cdot\nabla\mathcal{F}_{N}(0)\leqslant\frac{\mathcal{F}_{N}(ry)-\mathcal{F}_{N}(0)}{r}-y\cdot\nabla\mathcal{F}_{N}(0)+\frac{\delta_{N}(r)}{r}$
@@
~$\frac{\widetilde{\mathcal{F}}_{N}(0)-\widetilde{\mathcal{F}}_{N}(-ry)}{r}-y\cdot\nabla\mathcal{F}_{N}(0)\geqslant\frac{\mathcal{F}_{N}(0)-\mathcal{F}_{N}(-ry)}{r}-y\cdot\nabla\mathcal{F}_{N}(0)-\frac{\delta_{N}(r)}{r}.$
By the standard concentration result (e.g. see [20, Theorem 1.2]), there is a constant $C>0$ such that $\sup_{r\in(0,1]}\mathbb{E}\delta_{N}(r)\leqslant CN^{-\frac{1}{2}}$. This along with Lemma 2.1 and (A.1) gives

### v2

@@ 3. Approach via constrained free energy
We remark that, to our best knowledge, this proposition is new because the convexity of $\xi$ is only assumed to be over $\mathbf{S}^{D}_{+}$ as in (H4). Under a stronger assumption that $\xi$ is convex on $\mathbb{R}^{D\times D}$, this proposition is direct consequence of [44, Theorem 2 and Lemma 2]. The stronger convexity is needed in the proof of the upper bound [44, Lemma 2] to use Guerra’s interpolation.
@@
It is proven in [44, Theorem 2] that
@@
Since $\inf_{\pi:\>\pi(1)=z}\mathscr{P}^{\mathbf{s}}(\pi,y)\geqslant\inf_{\pi}\mathscr{P}^{\mathbf{s}}(\pi,y)=\mathscr{P}^{\mathbf{s}}(y)$, we get the lower bound for (3.6). The Parisi functional in [44, Theorem 2] is written in a different notation. It is explained in [17, Appendix A] how to rewrite it in our notation.
@@
where $\left\langle\cdot\right\rangle$ is the Gibbs measure naturally associated with $\mathcal{F}_{N}(0,y)$. Notice that $m_{N}\in\mathcal{K}$. Let $L$ be the Lipschitz coefficient of $\xi$ on $\mathcal{K}$. Then, we can use the definition of $\Sigma_{\varepsilon}(z)$ in (3.5) to see that the right-hand side in the display is bounded from above by $\frac{1}{2}L\varepsilon+|z|\varepsilon$. This along with Lemma 2.1 (with $h=\mathbf{s}$) implies that
@@
where $\left\langle\cdot\right\rangle$ is the Gibbs measure associated with $F^{0}_{N}\left(\Sigma_{\varepsilon}(z)\right)$. Due to the constraint imposed by $\Sigma_{\varepsilon}(z)$, we have $\left|m_{N}-z\right|\leqslant\varepsilon$ a.s. under $\left\langle\cdot\right\rangle$. Letting $C$ be the Lipschitz coefficient of $G$ on $\mathcal{K}$, we can obtain (3.7).
@@
The proof for the upper bound is contained in the proof of [44, Lemma 3]. For completeness, we present it below. Temporarily fix any $\delta>0$. By Proposition 3.2 and (3.7), for every $z\in\mathcal{K}$, there is $\varepsilon_{z}>0$ such that
@@
The standard Gaussian concentration inequalities (e.g. [40, Theorem 1.2]) gives a constant $C$ such that $\mathbb{E}\left|\widetilde{F}^{G}_{N}(S)-F^{G}_{N}(S)\right|\leqslant CN^{-\frac{1}{2}}$ uniformly in $N$ and $S$. Therefore, we get
@@ Appendix A Convergence of $m_{N}$
As mentioned in Remark 1.4, we can show that $m_{N}$ in (1.4) always converges under the Gibbs measure associated with $F^{\mathsf{soc}}_{N}$ in (1.2). When $h=\mathbf{s}$, we have that $m_{N}=\frac{\sigma\sigma^{\intercal}}{N}$ is the self-overlap and such a result has been proved in [17, Theorem 1.1 (1) and (2)]. A straightforward modification gives the desired result below.
@@
where $\mathscr{P}^{h}$ is defined in (2.4) and $\left\langle\cdot\right\rangle$ is the Gibbs measure associated with $F^{\mathsf{soc}}_{N}$ in (1.2).
@@
Let $\left\langle\cdot\right\rangle$ be associated with $F^{\mathsf{soc}}_{N}$. If $h$ is bounded and measurable, then
@@
Varying $y$, we get $\lim_{N\to\infty}\nabla\mathcal{F}_{N}(0,0)=\nabla\mathscr{P}^{h}(0)$ in $\mathbb{R}^{d}$. Recall from (2.11) that $\nabla\mathcal{F}_{N}(0,0)=\mathbb{E}\left\langle m_{N}\right\rangle$ where $\left\langle\cdot\right\rangle$ is associated with $\mathcal{F}_{N}(0,0)$. The desired result follows from the observation that $\mathcal{F}_{N}(0,0)=F^{\mathsf{soc}}_{N}$. ∎
@@
Let $\left\langle\cdot\right\rangle$ be associated with $F^{\mathsf{soc}}_{N}$. If $h$ is bounded and measurable, then
@@
Recall $\widetilde{\mathcal{F}}_{N}$ below (2.6) and we write $\widetilde{\mathcal{F}}_{N}(\cdot)=\widetilde{\mathcal{F}}_{N}(0,\cdot)$ for brevity. We can use (2.11) to rewrite
@@
In view of (2.12), $\widetilde{\mathcal{F}}_{N}$ is convex, which implies that $y\cdot\nabla\widetilde{\mathcal{F}}_{N}(0)-y\cdot\nabla\mathcal{F}_{N}(0)$ is bounded from above by
$\frac{\widetilde{\mathcal{F}}_{N}(ry)-\widetilde{\mathcal{F}}_{N}(0)}{r}-y\cdot\nabla\mathcal{F}_{N}(0)\leqslant\frac{\mathcal{F}_{N}(ry)-\mathcal{F}_{N}(0)}{r}-y\cdot\nabla\mathcal{F}_{N}(0)+\frac{\delta_{N}(r)}{r}$
@@
$\frac{\widetilde{\mathcal{F}}_{N}(0)-\widetilde{\mathcal{F}}_{N}(-ry)}{r}-y\cdot\nabla\mathcal{F}_{N}(0)\geqslant\frac{\mathcal{F}_{N}(0)-\mathcal{F}_{N}(-ry)}{r}-y\cdot\nabla\mathcal{F}_{N}(0)-\frac{\delta_{N}(r)}{r}.$
By the standard concentration result (e.g. see [40, Theorem 1.2]), there is a constant $C>0$ such that $\sup_{r\in(0,1]}\mathbb{E}\delta_{N}(r)\leqslant CN^{-\frac{1}{2}}$. This along with Lemma 2.1 and (A.1) gives
@@
Data availability. No datasets were generated during this work.
Conflict of interest. The author has no conflicts of interest to declare.